1. Schedule of cash payments for a service company

Horizon Financial Inc. was organized on February 28. Projected selling and administrative expenses for each of the first three months of operations are as follows:

March $172,200
April 158,400
May 144,100

Depreciation, insurance, and property taxes represent $36,000 of the estimated monthly expenses. The annual insurance premium was paid on February 28, and property taxes for the year will be paid in June. 74% of the remainder of the expenses are expected to be paid in the month in which they are incurred, with the balance to be paid in the following month.

Prepare a schedule of cash payments for selling and administrative expenses for March, April, and May.
Answer #1
Horizon Financial Inc.
Schedule of Cash Payments for Selling and Administrative Expenses
For the Three Months Ending May 31, 2016
March April May
March expenses:
Paid in March 100788
Paid in April 35412
April expenses:
Paid in April 90576
Paid in May 31824
May expenses:
Paid in May 79994
Total cash payments 100788 125988 111818
Workings:
March April May
March expenses:
Paid in March =(172200-36000)*74%
Paid in April =(172200-36000)*26%
April expenses:
Paid in April =(158400-36000)*74%
Paid in May =(158400-36000)*26%
May expenses:
Paid in May =(144100-36000)*74%